The rate constant of second order reaction is 5.7 x 10-5 dm3 mol-1
s
-1 at 298 K and 1.64 x
10-4 dm3 mol-1
s
-1 at 313 K. Calculate the activation energy and Arrhenius-pre
exponential factor.
1
Expert's answer
2020-12-17T07:03:45-0500


lnK= lnA- Ea/RT

Ln5.7×10-5=lnA-Ea/ (298×8.314)

2477.572 (Ln5.7×10-5) = 2477.572- Ea

- Ea=-24211.97-2477.572lnA…. (i)

Ln1.64×10-4= lnA-Ea/ (313×8.314)

2602.282(Ln1.64×10-4) = 2602.282lnA- Ea

- Ea=-22744.82-2602.282 lnA…. (ii)

-24211.97-2477.572lnA= -22744.82-2602.282 lnA

LnA= 11.76

A=31.979

- Ea=-24211.97-2477.572lnA

- Ea=-24211.97-2477.572(11.76)

Ea=53348.21672J/mol